Use simultaneous equations to solve:
Two numbers add up to 1. One number is 5 less than than the square of the other number. What are the numbers?

Answer by Boreal(15235)

x+y=1
x+5=y^2
so x=-y+1
x=y^2-5
y^2-5=-y+1
y^2+y-6=0
(y+3)(y-2)=0
y=-3 and 2
x=4 and -1
(4, -3)
(-1, 2)